OpenStudy (anonymous):

HELP!!! There are 5 times as many tomato plants in the garden as there are cucumber plants. The number of pepper plants is 7 more than the sum of the number of tomato and cucumber plants. There are 19 pepper plants in the garden. How many cucumber plants are there? How many tomato plants are there?

OpenStudy (anonymous):

let the number of cucumber plant be x then no. of tomato plant is 5x and the no of pepper plant is (x+5x)+7 by the given problem we have (x+5x)+7 = 19 now just solve the linear equation

OpenStudy (anonymous):

Would X represent 1?

OpenStudy (anonymous):

Because (1x + 5x) + 7 =13. not nineteen

OpenStudy (anonymous):

6x + 7 =19 6x =19-7 or 6x =12 or x=12/6 x=2 hence then no. of tomato plant is 5x =5*2=10

OpenStudy (anonymous):

So you use inverse operation to get the number of cucumber plants?

OpenStudy (anonymous):

So the number of Cucumber plants is 9, right?